How many milliliters of a 31% alcohol solution must be added to 6 milliliters of a 11% alcohol solution to produce a 21% alcohol solution?
Found 2 solutions by Alan3354, Maths68:
Answer by Alan3354(69443)

How many milliliters of a 31% alcohol solution must be added to 6 milliliters of a 11% alcohol solution to produce a 21% alcohol solution?
-----------
21% is the average, so it's equal amounts
6 ml

Answer by Maths68(1474)

Solution A
Amount = x ml
Concentration = 31% = 0.31

Solution B
Amount = 6 ml
Concentration = 11% = 0.11

Mixture
Amount = 6+x ml
Concentration = 21% = 0.21

(Amount of Solution A)*(Concentration of Solution A)+ (Amount of Solution B)*(Concentration of Solution B)=(Amount of Mixture)*(Concentration of Mixture)
(x)*(0.31)+ (6)*(0.11)=(6+x)*(0.21)
0.31x+0.66=1.26+0.21x
0.31x-0.21x=1.26-0.66
0.1x=0.6
0.1x/0.1=0.6/0.1
x=6


6ml of a 31% alcohol solution must be added to 6 ml of a 11% alcohol solution to produce 12ml of a 21% alcohol solution.